A jar contains 8 green, 4 blue, 10 red, and 2 yellow Skittles. A Skittle is randomly drawn, replaced, then another is drawn. Find each probability.


1
Expert's answer
2021-05-04T13:14:06-0400

The total number of skittles is

8+10+4+2=24

Drawing the skittles with replacement the items be independent that means that one event does not affect the outcome of the other

P(R) = 10/24, P(Y) =2/24, P(B) =4/24

P(G) =8/24

1. Probability of Red then yellow

P(R)P(Y) =10/24 *2/24

=5/144

2. Probability of blue then Green

P(B)P(G) =4/24*8/24

=1/8

3. Probability of Red then Blue

P(R)P(B) =10/24*4/24

=5/72